OpenStudy (anonymous):

Which is the solution? -4x + 4 - 2x = -2(3x - 2)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

-2 , 3/4, ?

OpenStudy (luigi0210):

Looking for ansurs?

OpenStudy (anonymous):

looking for help

OpenStudy (luigi0210):

Good

OpenStudy (skullpatrol):

$$\huge -4x + 4 - 2x = -2(3x - 2)$$

OpenStudy (luigi0210):

@skullpatrol 's got you then

OpenStudy (skullpatrol):

Let's start with what is the objective here? That is what is the question asking for?

OpenStudy (anonymous):

ok so now i got -4x + 4 -2x = 6 - -4

OpenStudy (anonymous):

solution

OpenStudy (skullpatrol):

Solution to what?

OpenStudy (anonymous):

the equation?

OpenStudy (skullpatrol):

Yes, what is a solution?

OpenStudy (anonymous):

i need you help lol thats why i asked this question im stuck after this point..

OpenStudy (skullpatrol):

Give me a definition of what is a "solution set" please.

OpenStudy (skullpatrol):

You need to be very clear about what you are doing in math and why you are doing it.

OpenStudy (skullpatrol):

Your question says "Which is the solution?" right??